OpenStudy (anonymous):

Line D passes through the points (1, 1) and (5, 6). What is the slope of a line perpendicular to Line D?

OpenStudy (isaidavila):

Fi D the first slope m=y2-y1/x2-x1 m=6-1/5-1 m=5/4 So when two lines are perpendicular the multiplication of their slopes is -1 m1*m2=-1 5/4*m2=-1 m2=-4/5

OpenStudy (anonymous):

thank you